# End-to-end Autonomous Driving: Challenges and Frontiers

### v2

@@ 1 Introduction
In this survey, we conduct an extensive review of this emerging topic. Fig. 1 provides an overview of our work. We begin by discussing the motivation and roadmap for end-to-end autonomous driving systems. End-to-end approaches can be broadly classified into imitation and reinforcement learning, and we give a brief review of these methodologies. We cover datasets and benchmarks for both closed- and open-loop evaluation. We summarize a series of critical challenges, including interpretability, generalization, world models, causal confusion, etc. We conclude by discussing future trends that we think should be embraced by the community to incorporate the latest developments from data engines, and large foundation models, amongst others. Note that this review is mainly orchestrated from a theoretical perspective. Engineering efforts such as version control, unit testing, data servers, data cleaning, software-hardware co-design, etc., play crucial roles in deploying the end-to-end technology. Publicly available information regarding the latest practices on these topics is limited. We invite the community towards more openness in future discussions.
Figure 1: Survey at A Glance. (a) Pipeline and Methods. We define end-to-end autonomous driving as a learning-based algorithm framework with raw sensor input and planning/control output. We deepdive into 270+ papers and categorize into imitation learning (IL) and reinforcement learning (RL). (b) Benchmarking. We group popular benchmarks into closed-loop and open-loop evaluation, respectively. We cover various aspects of closed-loop simulation and the limitations of open-loop evaluation for this problem. (c) Challenges. This is the main section of our work. We list key challenges from a wide range of topics and extensively analyze why these concerns are crucial. Promising resolutions to these challenges are covered as well. (e) Future Trends. We discuss how end-to-end paradigm could benefit by aid of the rapid development of foundation models, visual pre-training, etc. Partial photos by courtesy of online resources.
@@ 1.4 Contributions
(b) We extensively investigate the critical challenges that concurrent approaches face. Out of the more than 250 papers surveyed, we summarize major aspects and provide in-depth analysis, including topics on generalizability, language-guided learning, causal confusion, etc.
@@ 2.1.2 Inverse Optimal Control
Regarding cost design, it has representations including a learned cost volume in a bird’s-eye-view (BEV) [32], joint energy calculated from other agents’ future motion [69], or a set of probabilistic semantic occupancy or freespace layers [39, 70, 71]. On the other hand, trajectories are typically sampled from a fixed expert trajectory set [1] or processed by parameter sampling with a kinematic model [32, 39, 38, 70]. Then, a max-margin loss is adopted as in classic IOC methods to encourage the expert demonstration to have a minimal cost while others have high costs.
@@ 3.1 Real-world Evaluation
Early efforts on benchmarking self-driving involved real-world evaluation. Notably, DARPA initiated a series of races to advance autonomous driving. The first event offered \$1M in prize money for autonomously navigating a 240km route through the Mojave desert, which no team achieved [85]. The final series event, called the DARPA Urban Challenge, required vehicles to navigate a 96km mock-up town course, adhering to traffic laws and avoiding obstacles [86]. These races fostered important developments in autonomous driving, such as LiDAR sensors. Following this spirit, the University of Michigan established MCity [87], a large controlled real-world environment designed to facilitate testing autonomous vehicles. Besides such academic ventures, industries with the resources to deploy fleets of driverless vehicles also rely on real-world evaluation to benchmark improvements in their algorithms.
@@ 3.2.4 Vehicle Dynamics Simulation
CARLA
CoRL [90], noCrash [123], Town05 [6], LAV [52], Roach [21], Longest6 [28], Leaderboard v1 and v2 [13]
nuPlan
NAVSIM [124], Val14 [125], Leaderboard [14]
@@ 4.2.1 Representation Design
Naive representations are extracted with various backbones. Classic convolutional neural networks (CNNs) still dominate, with advantages in translation equivariance and high efficiency [175]. Depth-pre-trained CNNs [176] significantly boost perception and downstream performance. In contrast, Transformer-based feature extractors [177, 178] show great scalability in perception tasks while not being widely adopted for end-to-end driving yet. For driving-specific representations, researchers introduce the concept of bird’s-eye-view (BEV), fusing different sensor modalities and temporal information within a unified 3D space [179, 180, 132, 131]. It also facilitates easy adaptions to downstream tasks [2, 30, 181]. In addition, grid-based 3D occupancy is developed to capture irregular objects and used for collision avoidance in planning [182]. Nevertheless, the dense representation brings huge computation costs compared to BEV methods.
@@ 4.3 Complexity of World Modeling for Model-based RL
It is worth noting that learning world models in raw image space is non-trivial for autonomous driving. Important small details, such as traffic lights, would easily be missed in predicted images. To tackle this, GenAD [203] and DriveWM [204] employ the prevailing diffusion technique [205]. MILE [206] incorporates the Dreamer-style world model learning in the BEV segmentation space as an auxiliary task besides imitation learning. SEM2 [136] also extends the Dreamer structure but with BEV map inputs, and uses RL for training. Besides directly using the learned world model for MBRL, DeRL [195] combines a model-free actor-critic framework with the world model, by fusing self-assessments of the action or state from both models.
@@ 4.6 Lack of Interpretability
Rules Integration and Cost Learning: As discussed in Sec. 2.1.2, cost learning-based methods share similarities with traditional modular systems and thus exhibit a certain level of interpretability. NMP [32] and DSDNet [226] construct the cost volume in conjunction with detection and motion prediction results. P3 [39] combines predicted semantic occupancy maps with comfort and traffic rules constraints to construct the cost function. Various representations, such as probabilistic occupancy and temporal motion fields [1], emergent occupancy [71], and freespace [70], are employed to score sampled trajectories. In [227, 38, 125], human expertise and pre-defined rules including safety, comfort, traffic rules, and routes based on perception and prediction outputs are explicitly included to form the cost for trajectory scoring, demonstrating improved robustness and safety.
@@ 5.4 Foundation Model
Recent advancements in foundation models in both language [166, 167, 274] and vision [275, 273, 276] have proved that large-scale data and model capacity can unleash the immense potential of AI in high-level reasoning tasks. The paradigm of finetuning [277] or prompt learning [278], optimization in the form of self-supervised reconstruction [279] or contrastive pairs [165], etc., are all applicable to the end-to-end driving domain. However, we contend that the direct adoption of LLMs for driving might be tricky. The output of an autonomous agent requires steady and accurate measurements, whereas the generative output in language models aims to behave like humans, irrespective of its accuracy. A feasible solution to develop a “foundation” driving model is to train a world model that can forecast the reasonable future of the environment, either in 2D, 3D, or latent space. To perform well on downstream tasks like planning, the objective to be optimized for the model needs to be sophisticated enough, beyond frame-level perception.
@@ 6 Conclusion and Outlook
Outlook. The industry has dedicated considerable effort over the years to develop advanced modular-based systems capable of achieving autonomous driving on highways. However, these systems face significant challenges when confronted with complex scenarios, e.g., inner-city streets and intersections. Therefore, an increasing number of companies have started exploring end-to-end autonomous driving techniques specifically tailored for these environments. It is envisioned that with extensive high-quality data collection, large-scale model training, and the establishment of reliable benchmarks, the end-to-end approach will soon surpass modular stacks in terms of performance and effectiveness. In summary, end-to-end autonomous driving faces great opportunities and challenges simultaneously, with the ultimate goal of building generalist agents. In this era of emerging technologies, we hope this survey could serve as a starting point to shed new light on this domain.
@@ Acknowledgments
Li Chen, Penghao Wu, and Hongyang Li were partially supported by National Key R&D Program of China (2022ZD0160104), NSFC (62206172), and Shanghai Committee of Science and Technology (23YF1462000). Andreas Geiger and Bernhard Jaeger were supported by the ERC Starting Grant LEGO-3D (850533), the BMWi in the project KI Delta Learning (project number 19A19013O) and the DFG EXC number 2064/1 - project number 390727645. Kashyap Chitta was supported by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039A. We thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Bernhard Jaeger and Kashyap Chitta.

### v3

@@ 1 Introduction
In this survey, we conduct an extensive review of this emerging topic. Fig. 1 provides an overview of our work. We begin by discussing the motivation and roadmap for end-to-end autonomous driving systems. End-to-end approaches can be broadly classified into imitation and reinforcement learning, and we give a brief review of these methodologies. We cover datasets and benchmarks for both closed and open-loop evaluation. We summarize a series of critical challenges, including interpretability, generalization, world models, causal confusion, etc. We conclude by discussing future trends that we think should be embraced by the community to incorporate the latest developments from data engines, and large foundation models, amongst others. Note that this review is mainly orchestrated from a theoretical perspective. Engineering efforts such as version control, unit testing, data servers, data cleaning, software-hardware co-design, etc., play crucial roles in deploying the end-to-end technology. Publicly available information regarding the latest practices on these topics is limited. We invite the community towards more openness in future discussions.
Figure 1: Survey at A Glance. (a) Pipeline and Methods. We define end-to-end autonomous driving as a learning-based algorithm framework with raw sensor input and planning/control output. We deepdive into 270+ papers and categorize into imitation learning (IL) and reinforcement learning (RL). (b) Benchmarking. We group popular benchmarks into closed-loop and open-loop evaluation, respectively. We cover various aspects of closed-loop simulation and the limitations of open-loop evaluation for this problem. (c) Challenges. This is the main section of our work. We list key challenges from a wide range of topics and extensively analyze why these concerns are crucial. Promising resolutions to these challenges are covered as well. (d) Future Trends. We discuss how end-to-end paradigm could benefit by aid of the rapid development of foundation models, visual pre-training, etc. Partial photos by courtesy of online resources.
@@ 1.4 Contributions
(b) We extensively investigate the critical challenges that concurrent approaches face. Out of the more than 270 papers surveyed, we summarize major aspects and provide in-depth analysis, including topics on generalizability, language-guided learning, causal confusion, etc.
@@ 2.1.2 Inverse Optimal Control
Regarding cost design, it has representations including a learned cost volume in a bird’s-eye-view (BEV) [32], joint energy calculated from other agents’ future motion [69], or a set of probabilistic semantic occupancy or freespace layers [39, 70, 71]. On the other hand, trajectories are typically sampled from a fixed expert trajectory set [1, 72] or processed by parameter sampling with a kinematic model [32, 39, 38, 70]. Then, a max-margin loss is adopted as in classic IOC methods to encourage the expert demonstration to have a minimal cost while others have high costs.
@@ 3.1 Real-world Evaluation
Early efforts on benchmarking self-driving involved real-world evaluation. Notably, DARPA initiated a series of races. The first event offered \$1M in prize money for autonomously navigating a 240km route through the Mojave desert, which no team achieved [86]. The final series event, called the DARPA Urban Challenge, required vehicles to navigate a 96km mock-up town course, adhering to traffic laws and avoiding obstacles [87]. These races fostered important developments in self-driving, such as LiDAR sensors. Following this spirit, the University of Michigan established MCity [88], a large controlled real-world environment to facilitate testing autonomous vehicles. However, such academic ventures have not been widely employed for end-to-end systems due to a lack of data and vehicles. In contrast, industries with the resources to deploy fleets of driverless vehicles could rely on real-world evaluation to benchmark improvements in their algorithms.
@@ 3.2.4 Vehicle Dynamics Simulation
CARLA CoRL [91], noCrash [124], Town05 [6], LAV [52], Roach [21], Longest6 [28], Leaderboard v1 and v2 [13]
nuPlan NAVSIM [125], Val14 [126], Leaderboard [14]
@@ 4.2.1 Representation Design
Naive representations are extracted with various backbones. Classic convolutional neural networks (CNNs) still dominate, with advantages in translation equivariance and high efficiency [176]. Depth-pre-trained CNNs [177] significantly boost perception and downstream performance. In contrast, Transformer-based feature extractors [178, 179] show great scalability in perception tasks while not being widely adopted for end-to-end driving yet. For driving-specific representations, researchers introduce the concept of bird’s-eye-view (BEV), fusing different sensor modalities and temporal information within a unified 3D space [180, 181, 133]. It also facilitates easy adaptions to downstream tasks [2, 30, 182, 183]. In addition, grid-based 3D occupancy is developed to capture irregular objects and used for collision avoidance in planning [184]. Nevertheless, the dense representation brings huge computation costs compared to BEV methods.
@@ 4.3 Complexity of World Modeling for Model-based RL
It is worth noting that learning world models in raw image space is non-trivial for autonomous driving. Important small details, such as traffic lights, would easily be missed in predicted images. To tackle this, a few works [205, 206, 207] employ the prevailing diffusion technique [208]. MILE [209] incorporates the Dreamer-style world model learning in the BEV segmentation space as an auxiliary task besides imitation learning. SEM2 [137] also extends the Dreamer structure but with BEV map inputs, and uses RL for training. Besides directly using the learned world model for MBRL, DeRL [197] combines a model-free actor-critic framework with the world model, by fusing self-assessments of the action or state from both models.
@@ 4.6 Lack of Interpretability
Rules Integration and Cost Learning: As discussed in Sec. 2.1.2, cost learning-based methods share similarities with traditional modular systems and thus exhibit a certain level of interpretability. NMP [32] and DSDNet [229] construct the cost volume in conjunction with detection and motion prediction results. P3 [39] combines predicted semantic occupancy maps with comfort and traffic rules constraints to construct the cost function. Various representations, such as probabilistic occupancy and temporal motion fields [1], emergent occupancy [71], and freespace [70], are employed to score sampled trajectories. In [230, 38, 126, 183], human expertise and pre-defined rules including safety, comfort, traffic rules, and routes based on perception and prediction outputs are explicitly included to form the cost for trajectory scoring, demonstrating improved robustness and safety.
@@ 5.4 Foundation Model
Recent advancements in foundation models in both language [167, 168] and vision [276, 277] have proved that large-scale data and model capacity can unleash the immense potential of AI in high-level reasoning tasks. The paradigm of finetuning [278] or prompt learning [279], optimization in the form of self-supervised reconstruction [280] or contrastive pairs [166], etc., are all applicable to the end-to-end driving domain. However, we contend that the direct adoption of LLMs for driving might be tricky. The output of an autonomous agent requires steady and accurate measurements, whereas the generative output in language models aims to behave like humans, irrespective of its accuracy. A feasible solution to develop a “foundation” driving model is to train a world model that can forecast the reasonable future of the environment, either in 2D, 3D, or latent space. To perform well on downstream tasks like planning, the objective to be optimized for the model needs to be sophisticated enough, beyond frame-level perception.
@@ 6 Conclusion and Outlook
Outlook: The industry has dedicated considerable effort over the years to develop advanced modular-based systems capable of achieving self-driving on highways. However, these systems face significant challenges when confronted with complex scenarios, e.g., inner-city streets and intersections. Therefore, an increasing number of companies have started exploring end-to-end autonomous driving techniques specifically tailored for these environments. It is envisioned that with extensive high-quality data collection, large-scale model training, and the establishment of reliable benchmarks, the end-to-end approach will have enormous potential over modular stacks in terms of performance and effectiveness. In summary, end-to-end autonomous driving faces great opportunities and challenges simultaneously, with the ultimate goal of building generalist agents. In this era of emerging technologies, we hope this survey could serve as a starting point to shed new light on this domain.
@@ Acknowledgments
This project is partially supported by National Key R&D Program of China (2022ZD0160104), NSFC (62206172), and Shanghai Committee of Science and Technology (23YF1462000). Andreas Geiger and Bernhard Jaeger are supported by the ERC Starting Grant LEGO-3D (850533), the BMWi in the project KI Delta Learning (project number 19A19013O), and the DFG EXC number 2064/1 - project number 390727645. Kashyap Chitta is supported by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039A. We thank the International Max Planck Research School for Intelligent Systems for supporting Bernhard Jaeger and Kashyap Chitta.